OpenStudy (anonymous):

Solve (3/4)d = 3/8 Medal will be given!! Thank you in advance.(;

OpenStudy (imstuck):

Solving for d?

OpenStudy (anonymous):

Yes! lol

OpenStudy (zzr0ck3r):

multiply both sides by (4/3) and tell me what you get

OpenStudy (imstuck):

Actually I changed the fractions to decimals and found that way easier. Just sayin'.

OpenStudy (anonymous):

Lol what did you get? @IMStuck

OpenStudy (imstuck):

I did it like this: .75d=.375 d=.5 or 1/2

OpenStudy (zzr0ck3r):

dont change the fractions to decimals, that is missing the point.

OpenStudy (imstuck):

Very simple.

OpenStudy (zzr0ck3r):

because of course then how do you solve \(\frac{\pi}{\sqrt{2}}x=e\)

OpenStudy (imstuck):

You have a point, yes @zzr0ck3r But for the simplicity of simply finding d, you can't get any simpler.

OpenStudy (imstuck):

Listen to @zzr0ck3r so you can do the fraction stuff properly.

OpenStudy (zzr0ck3r):

but its missing the entire point.

OpenStudy (anonymous):

Lol I understand

OpenStudy (zzr0ck3r):

\(\frac{3}{4}d=\frac{3}{8}\) multiply both sides by \(\frac{4}{3}\) and you get \(\frac{4}{3}*\frac{3}{4}d=\frac{3}{8}*\frac{4}{3}\) \(\frac{\cancel{4}}{\cancel{3}}*\frac{\cancel{3}}{\cancel{4}}d=\frac{\cancel{3}}{8}*\frac{4}{\cancel{3}}\) so \(d=\frac{4}{8}=\frac{1}{2}\)

OpenStudy (anonymous):

Thanks both of you.(; @zzr0ck3r @IMStuck